Find the annual perecentage yield. A bank offers an APR of 2.3% compounded daily.
Answer by Theo(13342)

assuming the number of days in the year is 365, the calculations are done as follows:

annual percentage rate = 2.5%
nominal annual interest rate = apr / 100% = .025
number of compounding periods per year = 365
daily interest rate = .025 / 365 = .00006849315068
add 1 to that to get 1.00006849315068
raise that to the 365th power to get (1.00006849315068)^365 = 1.025314243
subtract 1 from that to get .025314243
that's the effective annual interest rate.
multiply that by 100% to get 2.5314243%
that's the annual percentage yield.

if you invest a dollar today and you get 2.5% annual percentage rate, you will have 1.025314243 dollars at the end of the year.

your effective interest rate is equal to .025314243
your annual percentage yield is 2.5314243%.